OpenStudy (anonymous):

How do I solve this? Determine the solution set of (x + 2)^2 = 36. a) {6, -8} b) {4, -8} c) {4, -6}

OpenStudy (anonymous):

What you need to remember is that a positive number and it's additive inverse both give you the same number when they are squared. Eg; 4 and -4 both give you 16 when they are squared. The square root of 36 is 6. The solutions for x then must be either 6 squared or -6 squared. x+2=6 and x+2=-6 then are equations that give the two solutions. x+2=6 x=4 x+2=-6 x=-8 (4,-8)